# ![[Uncaptioned image]](x1.png)

### v1

@@ 5 Discussion
Design choices. The main strength of DimVis lies in its model-agnostic approach and its flexible workflow. This independence from specific methods like UMAP or EBM ensures its adaptability to emerging algorithms. However, both choices were made deliberately since they are state-of-the-art methods that can scale to thousands of data entries (especially with DimVis’ zoom interaction) and hundreds of features without facing issues, such as how to visualize decision trees as non-space-filling node-link diagrams and simplify dimension examination to single or pairs of features.
Limitations and future work. DimVis currently only supports numeric features, limiting its use in scenarios involving categorical data. While it operates without class labels, the absence of these labels restricts EBM’s effectiveness in displaying an overview and global feature importances. Future work for DimVis includes extending support to non-numeric features and refining the user experience based on feedback from domain experts (as in Section 4). A formal user study is also planned to validate our findings further.

### v2

@@ 5 Discussion
Design choices. Although in its current version DimVis uses UMAP and EBM, the architecture of the tool allows it to be model-agnostic and flexible. The choices of EBM and UMAP were made since they are both state-of-the-art methods that can scale to large and high-dimensional datasets, but the flexible design of the workflow ensures its potential adaptability to emerging, future algorithms. Also, one of the strengths of EBM is that it can be interpreted with simple visualizations of single or pairs of features (as opposed to, for example, visualizing decision trees as non-space-filling node-link diagrams). The visualizations used for DimVis are inherited from the EBM model (pre-evaluated, for example, with real-world medical data [CLG∗15]) and do not represent a novelty of this paper. As such, a future direction could be to enhance the presentation beyond what is currently used. However, more complex visualizations may have a steeper learning curve and suffer in terms of scalability, effectiveness, and ease of use.
Limitations and future work. One important future work is an objective comparison to other similar tools and techniques so as to determine concretely how the detected explanations (e.g., features and their ranks) differ from other possible methods for generating similar explanations. This is one of our short-term plans for future work. We also intend to refine the user experience based on feedback from domain experts (as in Section 4) and perform a formal user study to validate our findings further and gather more feedback on possible improvements. Different possible visualization techniques (with variable degrees of complexity) will be tested with users to determine whether it is worth it to move beyond the current ones and how to do it. We intend to investigate ways to improve the computational efficiency of the underlying model by, for example, using GPU-based implementations of both the DR and the EBM algorithms. Finally, DimVis can currently only compare one data selection against the remaining projection, making a direct comparison between two subclusters a somewhat inefficient task. We intend to test the ability to select and compare two specific data subsets directly to streamline this process.